Astronomy Report – January 9th

Astronomy Report
Name: Agnieszka Pokrywka
Crew: 238
Date: 01/09/2022

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: Montana Learning Center
Objects Viewed: IC 1848, IC 1805, Ceres
Problems Encountered: I’m working now on the final images of all the observed objects using Astronomy Laptop. I encountered a problem with the calibration of the images though. Question: do I need to download calibration files for each observation separately or some observations will have the same files? I’m asking about that in order to save the bandwidth whenever possible.
I also wonder, is it possible to observe comets, like Leonard or Bernardinelli-Bornstein, with requested telescope? I can’t find them via the Skynet search.

MUSK OBSERVATORY
Not used.

Sol Summary – January 9th

Crew 238 Sol Summary Report 09Jan2022

Sol:7

Summary Title: Welcoming a Visitor to MDRS

Author’s name: Sionade Robinson

Mission Status: Ongoing

Sol Activity Summary: Today we welcomed Oliver Conroy, a feature
journalist with the Guardian newspaper, to MDRS. The visit had been
proposed by the Mars Society prior to our rotation, and accommodated
by the crew. After some final COVID testing by our HSO, we admitted
our visitor into the Habitat, shared our research work, gave him a
tour of the campus and entertained him to a lunch of lentil soup, bean
casserole and cinnamon apple crumble. He also greatly enjoyed
sourdough wholewheat bread cooked by Aga Pokrywka. He remarked he
hadn’t expected we would be eating so well. After lunch our visitor
accompanied three crew to walk the temporary labyrinth drawn yesterday
and witness an Emergency Drill (astronaut rescue) successfully
executed. The journalist left us at 4pm with an invitation to share
further crew experiences post rotation.

Look Ahead Plan: Two sequential afternoon EVA requests – the first a
visit to Area D for geologic, bacterial and radiation sampling and a
second short EVA to transport Dr Sandor to visit temporary labyrinth
drawn yesterday under her tuition and record walking experience.

Anomalies in work: None.

Weather: Sunny and warm. No wind

Crew Physical Status: Nominal.

EVA: Afternoon walking EVA (Robinson, Turner, Pokrywka) to walk
temporary labyrinth, followed by Emergency Drill exercise with two
person rescue team and prone astronaut recovery back to Hab.
Debriefed.

Reports to be filed: Sol, Operations, EVA Report, Journalist and HSO report.

Support Requested: Approval for two afternoon EVAs, one to Area D
(Pokrywka, Werner and Turner) for sampling and one to transport Dr
Sandor to record a walk in the temporary labyrinth drawn yesterday
under her tuition.

Additional Support: Collection and disposal of burnable and
non-burnable rubbish in the airlock. Thank you.

Operations Report – January 9th

Crew 238 Operations Report 09-01-2022

SOL:7

Name of person filing report: Simon Werner

Non-nominal systems: none

Notes on non-nominal systems: none

Spirit rover used: No

Hours: 149.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: No

Hours: 81.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 160.8

Beginning charge: 100%

Ending charge: 99%

Currently charging: yes

Perseverance rover used: No

Hours: 22.6

Beginning charge: unknown

Ending charge: 100%

Currently charging: Yes

General notes and comments: none

Summary of Hab operations:

WATER USE: 25 gallons

Water (static tank): 335

Water (loft tank): filled to 55 gallons

Water Meter: post pumping 01545921

Static to Loft Pump used – yes

Static tank pipe heater (on or off): on

Static tank heater (On or off) on

Toilet tank emptied: No

Summary of internet: nominal

Summary of suits and radios: nominal

Summary of GreenHab operations:

WATER USE: 8 gallons

Heater: On

Supplemental light: Yes

Harvest: no harvest today

Summary of Science Dome operations: The middle microscope was used for
optical analysis of bacterial culture by crew artist Pokryvka.

Dual split: Off

Summary of RAM operations: Checked battery state of a suit (Nr.4,
nominal) and placed 2 items there for pick up: a reset key and the
access code for the observatory’s door lock. (Observatory door is
locked). An instrument was found close to the science dome by the EVA
team today: description printed on it reads ”10 micron remote switch”.
I’ve put both items on the RAM working desk.

Summary of any observatory issues: none

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support:

There are two bags marked B and NB placed in the engineering airlock for pickup.

Thank You!